OpenStudy (anonymous):

which type of polynomials are not in 1 variable?????

OpenStudy (anonymous):

Variables usually refer to x y z for example, so x=5 would have 1 variable in it. If your question is asking about polynomials with more then one variable its probably refering to 3 space like z=x^2+y+3 for example

OpenStudy (anonymous):

im asking about which are ...like \[\frac{ 1 }{ x }\] and \[\sqrt{x}\] are the polynomials??????

jhonyy9 (jhonyy9):

so how do you think what will be the deifferents ? from ,,monom,binom and polinom "